OpenStudy (anonymous):

A store has 39 qt of milk some in punt cartons and some in quart cartons. There are six times as many quart cartons as pint cartons. How Many quart cartons are there

OpenStudy (anonymous):

Let's define the following to represent the number of pint and quart cartons, respectively.\[n_P, n_Q\]Now, we know that there are 2 pints per quart, so the total amount of milk is...\[\frac12n_P+n_Q = 39\]Now, we're told that the number of pint and quart cartons are related...\[n_Q=6n_P\]Let's put it all together now.\[\frac12n_P+6n_P=39\\\frac{13}2n_P=39\\13n_P=78\\n_P=6\]Now, we can determine the number of quart cartons.\[n_Q=6n_P=36\]